Refer to the following circuit (Q1 is a generic transistor), assuming β = 200 when the transistor is in the active region.

(1) If both R1 and R3 are short, calculate IB, IC, and IE.

(2) If R4 is short, calculate IB, VC, and IE.

(3) If both R2 and R3 are open, calculate IB, IC, and IE.

(4) If R1, R2 are swapped, calculate IB, IC, and IE.

R1 47ΚΩ R3 3.3kΩ 12V Vcc Q1 T R2 12kΩ R4 1.0kΩ
Answer #1

Ans 12V 1) 11c R3=0 (ie. Short) 1B •IE 12kas Ika 12 - 0.7 - Ile =0 le = 11.3 m IB. If 10.056 mA 11.3 201 Te + Ic - B18 = 11-23) rec IT 47kR M IB 12V LE 12-47163 - 0.7 (кл 201 IB = 0 11.3 240 = IB IB = 0.0455mA Ic BIB 9.112 mA If = (1+B) IR = 9-158mADear student hope it helps you. Please don't forget to give thumbs up thank you